### v1

@@ Abstract
We propose a novel visual place recognition approach, VOP, that efficiently addresses occlusions and complex scenes by shifting from traditional reliance on global image similarities and local features to image overlap prediction. The proposed method enables the identification of visible image sections without requiring expensive feature detection and matching. By focusing on obtaining patch-level embeddings by a Vision Transformer backbone and establishing patch-to-patch correspondences, our approach uses a voting mechanism to assess overlap scores for potential database images, thereby providing a nuanced image retrieval metric in challenging scenarios. VOP leads to more accurate relative pose estimation and localization results on the retrieved image pairs than state-of-the-art baselines on a number of large-scale, real-world datasets. The code is available at https://github.com/weitong8591/vop.
Keywords:
image retrieval image overlap relative camera pose vision transformer
@@ 1 Introduction
Image retrieval algorithms, particularly those concerning visual place recognition (VPR), play an essential role in computer vision and robotics [17, 15] problems. Place recognition is defined as the task of identifying the approximate location where a query image was taken, given a certain tolerance level [43, 77, 19, 2, 28, 85, 78, 24, 27, 36, 6, 31, 83]. This capability is crucial for estimating the state of robots and is widely utilized in autonomous, unmanned aerial, terrestrial, and underwater vehicles, as well as recent AR/VR devices.
@@
Typically, VPR is approached as an image retrieval problem, where the image to be localized is compared against a large database of posed images and, optionally, a 3D reconstruction of the scene. The most similar images retrieved from the database are used to estimate the precise location of the input query image, e.g., via local feature detection and matching. The complexity of VPR stems from a variety of factors such as perceptual aliasing, seasonal variations, changes in viewpoint, the presence of dynamic objects, illumination differences, occlusions, weather conditions, and intricate environments, as highlighted in the studies by Hong et al.[30], Doan et al.[20], and Subramaniam et al[72].
@@
Recent approaches conceptualize VPR as a task of estimating image similarity, conducting retrieval through learned image embeddings generated by a feature extraction architecture augmented with an aggregation or pooling mechanism, the most notable method being NetVLAD [2]. In this framework, images from the database are selected for retrieval based on their similarity to the query image, for example, by optimizing the cosine similarity between their embeddings. One notable drawback of this method is its failure to address partial overlaps resulting from occlusions. In practice, occlusions can significantly reduce image similarity scores, despite the unobstructed portions of the scene potentially offering valuable cues for localization.
@@
To be robust to occlusions and to provide a more nuanced understanding of the query image in relation to the database, we propose a novel approach for VPR. Instead of relying on global similarities or local features, we frame the problem as an overlap prediction, enabling us to methodically understand which parts of a particular image are visible without having to perform feature detection and matching. We approach the problem by generating embeddings for individual image patches rather than the entire image, thereby facilitating the efficient establishment of patch-to-patch correspondences. These matches are subsequently utilized within a voting mechanism to calculate an overlap score for potential database images, providing a nuanced metric for image retrieval amidst complex visual environments.
@@
Illustrative examples are provided in Fig. 1. In these examples, AnyLoc [33], a state-of-the-art retrieval solution, processes a query image alongside two database images: one sharing the scene with the query albeit with occlusion and another from a different environment. AnyLoc erroneously attributes higher similarity to the unrelated image. In contrast, our proposed approach successfully identifies the correct image and the areas likely to overlap.
Figure 1: An outdoor and indoor image retrieval example of cases when image similarity prediction by the state-of-the-art AnyLoc [33] methods fails, while the proposed overlap prediction selects the correct database image, as indicated by the retrieval scores visualized above/below the database images. The middle column shows the same scene as the query (left) with heavy occlusion. The right column depicts a different location.
The contributions of this paper are summarized as follows:
@@
    A novel approach, transitioning from global image similarity to overlap assessment which exhibits enhanced robustness to challenges such as occlusions. The proposed approach is built on patch-level analysis by learning patch embeddings suitable for visual localization. It implements a single radius search across embeddings of image patches in the database.
@@
    The introduction of a robust voting mechanism for image retrieval which demonstrates improvement over conventional similarity-based methods.
Moreover, we identify a limitation inherent in the standard metric, namely recall, for evaluating image retrieval accuracy as done in recent studies [5, 33]. Recall quantifies the success of a method in retrieving correct database images without considering any application-specific evaluation metrics. Consequently, we opt for the Area Under the Curve (AUC) metric, specifically focusing on relative pose accuracy, to provide a more nuanced and relevant performance assessment.
@@ 2 Related Work
Image retrieval methods for VPR can be categorized into two groups, one globally representing the whole image and the other locally for keypoints.
Localization the process of determining the position and orientation of an agent within a pre-built map, is pivotal across various domains such as autonomous driving [29], drones [40], and augmented reality [13, 45]. The differentiation in localization techniques arises from their scene representation methods – be it through explicit 3D models [25, 26, 32, 38, 42, 46, 60, 63, 64, 66, 73, 86, 62], sets of posed images [87, 88, 56, 7], or implicitly via neural network weights [35, 79, 81, 34, 8, 9, 4, 14, 11, 47] – and their approach to camera pose estimation, whether by 2D-3D [25, 26, 60, 62, 64, 73, 86] or 2D-2D [88, 7] matches, or through a composite of base poses [35, 81, 34, 65, 56, 47]. Enhanced robustness is often sought through sequential image analysis rather than single-frame methods, with [50] adapting sequence matching to graph search and [80] developing an online sequence matching technique for real-time localization. In practice, localization comprises two main steps: place recognition and precise localization stages. Here, we focus on the first step, finding potential locations of a query image.
@@
For measuring closeness, traditional methods like Bag of Visual Words [16] (BoW) identify key local image patches and store them as a codebook. BoW is often used in conjunction with 128-dimensional SIFT [44] descriptors both for map creation and image retrieval. These high-dimensional features often require approximate nearest neighbor (NN) search or GPUs to accelerate the search process. Showcasing the unbroken popularity of such approaches, they are still widely employed in applications, e.g., in ORB-SLAM [48, 49, 12].
As demonstrated by Zheng et al.[89], using CNNs for outperforming hand-crafted features in various applications, namely through compact (fixed-length) representations, has gradually become the norm. Most recent methods utilize learned embeddings generated by a feature extraction backbone with an aggregation or pooling head, the most notable method being NetVLAD [2]. In recent works, transformer-based retrieval methods [21] have been introduced, incorporating metric learning and attention mechanisms. These methods utilize local and global features, or a fusion of the two, to identify similar images.
Global retrieval methods integrate or combine handcrafted or learning-based local descriptors, such as SIFT [44] through the bag-of-visual-words model [55, 70], or employ alternative algorithms [76]. There is a growing trend towards learning-based techniques that devise improved methods for feature aggregation. NetVLAD [2] introduces an end-to-end trainable image representation extractor, employing weak supervision through soft assignment to cluster centers. Generalized Mean Pooling (GeM) [57] improves by introducing a learnable pooling layer that generalizes both max and average pooling, with a CNN serving as the backbone. Additionally, AP-GeM [59] enhances ranking accuracy by minimizing a relaxed global mean Average Precision (mAP) over a set of images. A CNN-based solution, CosPlace [5], utilizes the same CNN backbone and GeM pooling but approaches the place retrieval issue as a classification problem. MixVPR [1] transitions feature maps into a compact representation space through cascade MLP blocks. More recently, AnyLoc [33] emerges as a versatile method suitable for various image retrieval contexts, exploring CNN-based or Vision Transformer (ViT)-based foundational models and aggregation techniques to construct a vocabulary for searching in a large database.
@@ 3.1 Problem Statement
In place recognition problems, the aim is to find images $I_{j}\in\mathbb{R}^{H\times W\times C}$ containing the same landmark or scene as the query $I_{i}\in\mathbb{R}^{H\times W\times C}$ and distinguishable from others. Usually, it is done by transforming the images to a high-dimensional embedding space and measuring the distances to select the top-$k$ most similar images. In this embedding space, an image $I_{i}$ is represented by a high-dimensional global $\mathbf{g}_{i}\in\mathbb{R}^{d_{g}}$ or local descriptors $\mathbf{d}_{i},\in\mathbb{R}^{n\times d_{l}}$, using encoder $f(\cdot)$, where $n\in\mathbb{N}$ denotes the number of local features, and $d_{g}\in\mathbb{N}$ is set to one for the whole image, $d_{l}\in\mathbb{N}$ to 256, 128, etc. The matching score, then, can be computed as the similarities of the descriptors, $s(I_{i},I_{j})=sim(\mathbf{d}_{i},\mathbf{d}_{j})$ or $sim(\mathbf{g}_{i},\mathbf{g}_{j})$, for example, as their cosine similarity.
Instead of relying on a single score that quantifies the similarity of two images, we focus on patch-level representations and formulate the matching score as the overlap between the images. Let us split image $I$ into a set $\mathcal{P}$ of rectangular patches by a uniform grid consisting of $n\times n$ rows and columns. Given sets of such rectangular patches $\mathcal{P}_{i}$ and $\mathcal{P}_{j}$ in the two images, we can define a patch-level overlap score as follows:
@@
for each pair $(p,q)$, where $p\in\mathcal{P}_{i}$ and $q\in\mathcal{P}_{j}$ are the image patches, $\mathbf{R}\in\text{SO}(3)$ are the absolute rotations of the cameras, $\mathbf{t}\in\mathbb{R}^{3}$ are their absolute positions, $\mathbf{K}\in\mathbb{R}^{3\times 3}$ represents the intrinsic camera parameters, function $\text{cone}(\cdot)$ represents the 3D cone defined by the camera pose and the current patch, function $\text{volume}(\cdot,\cdot)$ measures the overlap of two posed cones in three dimensions. While this measure is easy to calculate given the camera poses and intrinsics, we observed that it does not reflect real-world applications, e.g., 3D reconstruction. Thus, we reformulate this measure as follows:
$o^{\prime}(p,q)=\sum_{X\in\mathcal{X}}\llbracket\text{inside}(\mathbf{K}_{i}(\mathbf{R}_{i}\mathbf{X}+\mathbf{t}_{i}),p)\rrbracket*\llbracket\text{inside}(\mathbf{K}_{j}(\mathbf{R}_{j}\mathbf{X}+\mathbf{t}_{j}),q)\rrbracket,$ (2)
@@
The overlap of two entire images $I_{i}$ and $I_{j}$ is defined as the sum of overlap scores across match pairs as follows:
$O(I_{i},I_{j})=\sum_{p\in\mathcal{P}_{i}}\sum_{q\in\mathcal{P}_{j}}o^{\prime}(p,q)\approx\sum_{(p,q)\in\mathcal{C}}o^{\prime}(p,q),$ (3)
where set $\mathcal{C}=\{(p,q)\;|\;p\in\mathcal{P}_{i},q=\arg_{w}\max o^{\prime}(p,w)\in\mathcal{P}_{j},w\in\mathcal{P}_{j}\}$ is a set of patch-to-patch correspondences defined by assigning the most overlapping patch from the database image to each patch in the query image. Please note that while we could easily define image overlap directly by Eq. 2, such a formulation does not allow for the next steps that allow learning an overlap score.
However, this formulation still does not allow for retrieval, as in practice, the camera pose corresponding to the query image is unknown, and our aim is to estimate it. To do so, we further approximate the image overlap as follows:
$O(I_{i},I_{j})\approx\sum_{(p,q)\in\mathcal{C}}\rho(sim(p,q)),$ (4)
where $\rho:\mathbb{R}\to\mathbb{R}$ is a robust function. For example, $\rho$ can be defined as a voting mechanism as $\rho(x)=\llbracket x>\epsilon\rrbracket$, where $\epsilon$ is a manually set threshold and $x$ is the patch similarity $sim(p,q)$.
@@
$I^{*}=\arg_{j\in 1,\dots,N}\max O(I_{i},I_{j}).$ (5)
In the scenario where multiple images need to be retrieved for a single query, we straightforwardly define a ranking by sorting images with high visual overlap. Thus, image retrieval from image collections becomes a patch-matching problem. Example matches are visualized in Fig. 2.
Figure 2: Patch matching examples in an image pair (column) with varying patch sizes (16, 64, and 256). When using 256 patches, the embeddings are directly provided by Dino V2 [51]. We used average pooling (16 and 64 patches) to calculate other resolutions. The numbers inside a particular patch denote the match index. The color overlay is calculated by applying PCA to the patch embeddings.
@@ 3.2 Learning Patch Matching via Patch Similarities
In order to define the similarity of image patches $p$ and $q$ in images $I_{i}$ and $I_{j}$, respectively, we aim to learn an embedding that allows for casting the problem as a descriptor matching. Thus, we aim to distill embeddings $e_{p}\in\mathbb{R}^{d}$ and $e_{q}\in\mathbb{R}^{d}$ such that $sim(p,q)=\delta(e_{p},e_{q})$ is high if and only if $p$ and $q$ are overlapping and zero otherwise. The function $\delta(\cdot)$ measures the similarity in the embedding space, e.g., as the cosine similarity.
Figure 3: The learning pipeline of the proposed Visual Overlap Prediction (VOP) method is shown, with a frozen ViT backbone [51], trainable encoder head, attention layers, and contrastive loss on the similarities. The input to the training is image pairs that are broken down into rectangular patches to learn patch-to-patch overlap scores.
Figure 4: The proposed Visual Overal Prediction (VOP) method at inference. Given an input query image and an image collection representing the reference map, the images are broken down into rectangular patches, the frozen backbone [51] extracts patch-level features, which are then fed into an encoder network, obtaining the final embeddings. The image retrieval is performed by running a radius search in the embedding space for each query patch to find potential database matches. The final overlap scores are determined by robust voting on the formed “query patch”-to-“database patch” correspondences. In practice, the map embeddings are obtained in an offline.
Pipeline for learning embeddings is visualized in Fig. 3. Given a pair of images $I_{i}$ and $I_{j}$, the problem is to learn embeddings $e_{p}$ and $e_{q}$ as defined previously. First, we obtain patch features by inputting the images into a Visual Transformer architecture. We use a frozen DINO V2 [51] backbone. These features are then fed into an encoder head, comprising one linear projection layer plus a few fully connected layers with GELU activation and dropout layer to encode the centers of the patches. Given the corresponding patch pair $(p,q)$ as a positive sample, we train this encoder to predict similar embeddings for $p$ and $q$. To do so, we employ self and cross-attention layers and contrastive learning on the similarities.
@@
For a query patch $p$, a set of candidate patches $\{q,q_{1},q_{2},...,q_{n}\}$ are provided for training, where $\{q_{1},q_{2},...,q_{n}\}$ act as $n$ negative samples, depicting patches that do not have a common field-of-view with the query patch. In our training pipeline, half of the negative samples come from images that overlap with the query image in general, but the selected patches do not. There are 30-90% negative patches in those generally overlapping image pairs. The remaining set of samples contains images capturing other scenes or other parts of the same scene.
We train our model by optimizing both a contrastive loss and an attention-based one as follows:
$\mathcal{L}=\alpha_{1}*\mathcal{L}_{\text{contrastive}}+\alpha_{2}*\mathcal{L}_{\text{attention}}.$ (6)
The contrastive loss is defined as follows:
@@
$\mathcal{L}_{\text{contrastive}}=\frac{\sum_{i}^{n}\sum_{j}^{n}l_{\text{GT}}(\sigma-sim(e_{p},e_{q}))^{2}-\sum_{i}^{n}\sum_{j}^{n}(1-l_{\text{GT}})sim(e_{p},e_{q})^{2}}{n^{2}},$ (7)
@@
where $l_{\text{GT}}$ is the ground-truth labels for each pair of patches, e.g., as the similarities between two patches from a image pair, $\llbracket sim(\mathbf{p}_{n_{i}},\mathbf{p}_{n_{j}})>\epsilon\rrbracket$, where $sim(\cdot)$ is the similarity measurement of the $d$-dimensional descriptors, i.e., cosine similarity with $\sigma=1$ as the margin. It penalizes the cases where there are high similarities on the patches but non-overlapping in the ground truth, i.e., no visible 3D points. We resize the input images to a fixed size. Therefore, the overall loss is averaged over $n^{2}$ tentative matching patches on one image pair.
Additionally, the self and cross attention layers $A(\cdot)$ are applied to select the relevant information and measure the similarity of two patch descriptors. Then we replace $sim(\cdot)$ by $A(\cdot)$ in Eq. 7, and the final loss, is computed by Eq. 6. The attentions are only used for training and do not change the descriptors in inference. Note that the same losses can be employed to project global descriptors, and the zero-overlapping image pairs can be used as the negative examples.
@@
In order to provide means for quick prefiltering, we add a classification (CLS) token to the employed Transformer network that is trained to distill global image embeddings based on the image overlap scores, shown in Table 5. These tokens will be used to select a subset of potentially overlapping images given the query.
@@
Constructing Training Data. Now, we introduce the visual supervision used to build GT labels for training. The ground-truth overlapping patches of the training image pairs are achieved following LightGlue [41] employing pixel-wise depth and relative pose. The ground-truth matching local patches are found in two steps: First, a set of dense pixel coordinates $c_{i}$ are detected in the first image $I_{i}$, and transformed to 3D points $X_{i}$. The obtained 3D point cloud is filtered by non-zero depths and transformed to the reference camera, i.e., $kp_{i2j}$, by the given transformation matrix. Given camera intrinsics, we project the 3D points $X_{i}$ to the camera planes and check for visibility. We make sure that only circle-consistent points are kept by using the projected point, calculating its 3D coordinates, and back-projecting to the original image to perform a visibility check again. The ground-truth matches are supposed to contain the points with a consistent depth [41]. This procedure provides the initial set of positive and negative pairs of matching image patches.
@@
Next, given the visible pixel matches, we first check which patches they belong to and define the matching patches by having at least one visible pixel correspondence. We assume one-to-one matching relations, i.e., the patch with the most patches matches in the image pair will be chosen if there are multiple. Note that the visual supervision is built mutually from the image to the reference and vice versa.

### v2

@@ Abstract
Visual place recognition methods struggle with occlusions and partial visual overlaps. We propose a novel visual place recognition approach based on overlap prediction, called VOP, shifting from traditional reliance on global image similarities and local features to image overlap prediction. VOP proceeds co-visible image sections by obtaining patch-level embeddings using a Vision Transformer backbone and establishing patch-to-patch correspondences without requiring expensive feature detection and matching. Our approach uses a voting mechanism to assess overlap scores for potential database images. It provides a nuanced image retrieval metric in challenging scenarios. Experimental results show that VOP leads to more accurate relative pose estimation and localization results on the retrieved image pairs than state-of-the-art baselines on a number of large-scale, real-world indoor and outdoor benchmarks. The code is available at https://github.com/weitong8591/vop.git.
@@ 1 Introduction
Visual Place Recognition (VPR) is defined as the task of identifying the approximate location where a query image was taken, given a certain tolerance level [30, 55, 14, 2, 20, 60, 56, 19, 21, 26, 7, 23, 58]. VPR methods have been widely used in computer vision and robotics [12, 10] problems, such as autonomous, unmanned aerial [32], terrestrial, and underwater vehicles [43], as well as recent AR/VR devices.
@@
Typically, VPR is approached as an image retrieval problem, where the image to be localized is compared against a large database of posed images and, optionally, a 3D reconstruction of the scene. Then, the most similar images retrieved from the database are used to estimate the precise location of the input query image, e.g., via local feature detection and matching. The complexity of VPR stems from a variety of factors such as seasonal variations, changes in viewpoint, the presence of dynamic objects, illumination differences, occlusions, weather conditions, and large-scale environments, as discussed in the studies by Hong et al.[22], Doan et al.[15], and Subramaniam et al[50].
Figure 1: An example where the SOTA AnyLoc [25] scores a negative DB image (right, at a different location) higher than an occluded positive example (left, the same scene as the query with heavy occlusion). VOP ranks the database (DB) images correctly.
@@
Recent approaches implement VPR as an estimation of image similarity and retrieval through learned image embeddings generated by a feature extraction architecture augmented with an aggregation or pooling mechanism, such as NetVLAD [2], CosPlace [6], and AnyLoc [25]. In such frameworks, images from the database are retrieved by optimizing their similarity between the embeddings to the query image, e.g., cosine similarity. One notable drawback of these methods is their sensitivity to partial overlaps resulting from occlusions. In practice, occlusions can significantly reduce image similarity scores despite the unobstructed portions of the scene, potentially offering valuable cues for localization. Second, such methods optimize the similarity of the embeddings on positives closer than the negative samples, which is hard to train as there is no ground truth (GT) prior to indicating the exact degree of similarity. Instead, we build GT patch matches as supervised by the 3D reconstruction. In addition, several reranking-based VPR methods [20, 62] retrieve the most similar images from the database on the shortlist built by global embeddings. Those improves the performance of retrieving correct images but needs more storage .
@@
Instead of relying on global similarities or local features, we frame the problem as overlap prediction. This enables us to methodically understand which parts of a particular image are visible without having to perform feature detection and matching. Also, focusing on overlaps makes VOP more robust to occlusions and provides a better understanding of the query image in relation to the database. VOP encodes individual patches rather than the entire image, thereby facilitating the efficient establishment of patch-to-patch correspondences. These matches are subsequently utilized within a voting mechanism to calculate an overlap score for potential database images, providing a nuanced metric for retrieval amidst complex visual environments.
@@
In Fig. 1, we show two retrieval examples using the proposed Visual Place Recognition by Overlap Prediction (VOP) and AnyLoc [25], a state-of-the-art retrieval solution. They process a query image alongside two database images: one sharing the scene with the query albeit with occlusion, and another from a different environment. AnyLoc erroneously attributes higher similarity to the unrelated image. In contrast, our proposed approach successfully identifies the correct image and the areas likely to overlap.
The contributions of this paper are as follows:
@@
    A novel approach focusing on patch overlap assessment, exhibiting enhanced robustness to challenges such as occlusions. The proposed method, VOP, learns patch embeddings suitable for visual localization and implements a single radius search across image patches in the database.
@@
    The introduction of a robust voting mechanism for image retrieval shows improvement over conventional similarity-based methods with nearest neighbors.
  3. 3.
    Breaking the frame of looking for close database images in standard VPR methods by considering geometric applications and using specific evaluation metrics.
@@ 2 Related Work
Image retrieval methods for VPR problems can be categorized into two groups, one globally representing the whole image and the other locally for keypoints using handcrafted or learning-based embeddings.
@@
To find the closest matches in the indexing structures, traditional methods like Bag of Visual Words [11] (BoW) identify key local image patches and store them as a codebook. BoW is often used in conjunction with 128-dimensional SIFT [31] descriptors both for map creation and image retrieval. These high-dimensional features often require approximate nearest neighbor (NN) search or GPUs to accelerate the searching process. Showcasing the unbroken popularity of such approaches, they are still widely employed in applications, e.g., in ORB-SLAM [33, 34, 9].
As demonstrated by Zheng et al.[61], using CNNs to improve hand-crafted features in various applications, namely through compact (fixed-length) representations, has gradually become the norm. Most recent methods utilize learned embeddings generated by a feature extraction backbone with an aggregation or pooling head, e.g., NetVLAD [2]. In recent works, transformer-based retrieval methods [16] have been introduced, incorporating metric learning and attention mechanisms. These methods utilize local and global features, or a fusion of the two, to identify similar images.
Global retrieval methods usually integrate or combine handcrafted or learning-based local descriptors, such as SIFT [31] through the BOW model [39, 48], or employ alternative algorithms [54]. There is a growing trend towards learning-based techniques that devise improved methods for feature aggregation. NetVLAD [2] introduces an end-to-end trainable image representation extractor, employing weak supervision through soft assignment to cluster centers. Generalized Mean Pooling (GeM) [40] proposes a learnable pooling layer that generalizes both max and average pooling, with a CNN serving as the backbone. Additionally, AP-GeM [42] enhances ranking accuracy by minimizing a relaxed global mean Average Precision (mAP) over a set of images. A CNN-based solution, CosPlace [6], utilizes the CNN backbone and GeM pooling but approaches the place retrieval issue as a classification problem. MixVPR [1] transitions feature maps into a compact representation space through cascade MLP blocks. AnyLoc [25] emerges as a versatile method suitable for various image retrieval contexts, exploring CNN-based or Vision Transformer (ViT)-based foundational models and aggregation techniques to construct a vocabulary for searching in a large database. More recently, DINOv2-SALAD [24] finetunes DINOv2 and proposes a new aggregation method based on optimal transport and the relations between features and clusters.
@@ 3.1 Problem Statement
In place recognition problems, the aim is to find images $I_{j}\in\mathbb{R}^{H\times W\times C}$ containing the same landmark or scene as the query $I_{i}\in\mathbb{R}^{H\times W\times C}$. Usually, it is done by transforming the images to a high-dimensional embedding space and measuring the distances to select the top-$k$ most similar images. In this embedding space, an image $I_{i}$ is represented by a high-dimensional global $\mathbf{g}_{i}\in\mathbb{R}^{d_{g}}$ or local descriptors $\mathbf{d}_{i},\in\mathbb{R}^{n_{l}\times d_{l}}$, using encoder $f(\cdot)$, where $n_{l}\in\mathbb{N}$ denotes the number of local features, with the dimensions $d_{g}\in\mathbb{N}$, $d_{l}\in\mathbb{N}$ The matching score, then, can be computed as the similarities of the descriptors, $s(I_{i},I_{j})=sim(\mathbf{d}_{i},\mathbf{d}_{j})$ or $sim(\mathbf{g}_{i},\mathbf{g}_{j})$, for example, as their cosine similarity.
Local features are usually image-dependent and redundant with variable lengths. Instead, we focus on patch-level representations and redefine the matching score as the overlap between the images. It is simple and efficient to achieve patch-level descriptors with less dependency. Let us split image $I$ into a set $\mathcal{P}$ of rectangular patches by a uniform grid consisting of $n$ rows and columns. Given sets of such rectangular patches $\mathcal{P}_{i}$ and $\mathcal{P}_{j}$ in the two images, there are multiple ways to define overlap. Here we define a patch-level overlap score as follows:
@@
for each pair $(p,q)$, where $p\in\mathcal{P}_{i}$ and $q\in\mathcal{P}_{j}$ are the image patches, $\mathbf{R}\in\text{SO}(3)$ are the absolute rotations of the cameras, $\mathbf{t}\in\mathbb{R}^{3}$ are absolute positions, $\mathbf{K}\in\mathbb{R}^{3\times 3}$ represents the intrinsic camera parameters, function $\text{cone}(\cdot)$ is the 3D cone defined by the camera pose and the current patch, function $\text{volume}(\cdot,\cdot)$ measures the overlap of two posed cones in three dimensions. While this measure is easy to calculate given the camera poses and intrinsics, we observed that it does not reflect real-world applications, e.g., 3D reconstruction. Thus, we reformulate this measure as follows:
Figure 2: Patch matching examples in an image pair (column) with varying patch sizes (16, 64, and 256). The 256 patches from DINOv2 [35] are average pooled to 16 and 64 patches. In each column, the numbers inside the patches indicate which ones are matched. The color overlay is calculated by applying PCA to the patch embeddings.
$\begin{split}o^{\prime}(p,q)=\sum_{X\in\mathcal{X}}\llbracket\text{inside}(\mathbf{K}_{i}(\mathbf{R}_{i}\mathbf{X}+\mathbf{t}_{i}),p)\rrbracket*\\ \llbracket\text{inside}(\mathbf{K}_{j}(\mathbf{R}_{j}\mathbf{X}+\mathbf{t}_{j}),q)\rrbracket,\end{split}$ (2)
@@
Eq. 2 defines the patch overlap as the number of 3D points commonly visible on both patches. Then the overlap of two entire images $I_{i}$ and $I_{j}$ is defined as the sum of overlap scores across matching patches as follows:
$O(I_{i},I_{j})=\sum_{p\in\mathcal{P}_{i}}\sum_{q\in\mathcal{P}_{j}}o^{\prime}(p,q)=\sum_{(p,q)\in\mathcal{C}}o^{\prime}(p,q),$ (3)
where set $\mathcal{C}=\{(p,q)\;|\;p\in\mathcal{P}_{i},q=\arg_{w}\max o^{\prime}(p,w)\in\mathcal{P}_{j},w\in\mathcal{P}_{j}\}$ is a set of patch-to-patch correspondences defined by assigning the most similar patch from the database image to each patch in the query.
However, this formulation still does not allow for retrieval, as the camera pose corresponding to the query image is unknown in practice. Thus, we further approximate the image overlap as the similarities between the patch embeddings in a robust way as follows:
$\hat{O}(I_{i},I_{j})=\sum_{(p,q)\in\mathcal{C}}\rho(sim(f(p),f(q))),$ (4)
where $f(\cdot)$ is an encoder with a single patch as input, usually using deep networks or hand-crafted methods, and $\rho:\mathbb{R}\to\mathbb{R}$ is a robust function, e.g., $\rho$ can be defined as a voting mechanism as $\rho(x)=\llbracket x>\epsilon\rrbracket$, where $\epsilon$ is a manually set threshold with $x$ as the patch similarity $sim(f(p),f(q))$.
@@
$I^{*}=\arg_{j\in 1,\dots,N}\max\hat{O}(I_{i},I_{j}).$ (5)
In the scenario where multiple images need to be retrieved for a single query, we straightforwardly define a ranking by sorting images with high visual overlap. Thus, image retrieval from collections becomes a patch-matching problem. Example matches are visualized in Fig. 2. Next, we will talk about how to learn the encoder $f(\cdot)$ in VOP focusing on accurate overlap prediction.
@@ 3.2 Learning Patch Matching via Patch Similarities
In order to find the similarity of image patches $p$ and $q$ in images $I_{i}$ and $I_{j}$, respectively, we aim to learn an embedding $f(\cdot)$ that allows for casting the problem as a descriptor matching. Thus, we aim to distill embeddings $e_{p}\in\mathbb{R}^{d}$ and $e_{q}\in\mathbb{R}^{d}$ such that $sim(p,q)(=\delta(e_{p},e_{q})=\delta(f(p),f(q))$) is high if and only if $p$ and $q$ are overlapping and zero otherwise. The function $\delta(\cdot)$ measures the similarity in the embedding space, e.g., as the cosine similarity.
Figure 3: The learning pipeline of the proposed Visual Overal Prediction (VOP) method is shown, with a frozen ViT backbone [35], trainable encoder head, and contrastive loss on the similarities. The input to the training is image pairs that are broken down into rectangular patches to learn patch-to-patch overlap scores.
Figure 4: The proposed Visual Overlap Prediction (VOP) method at inference. Given an input query image and an image collection representing the reference map, the images are broken down into rectangular patches, the frozen backbone [35] extracts patch-level features, which are then fed into an encoder network, obtaining the final embeddings. The image retrieval is performed by running a radius search in the embedding space for each query patch to find potential database matches. The final overlap scores are determined by robust voting on the formed “query patch”-to-“database patch” correspondences. In practice, the map embeddings are obtained in an offline.
The pipeline for learning paytch-level embeddings is visualized in Fig. 3. Given a pair of images $I_{i}$ and $I_{j}$, the problem is to learn embeddings $e_{p}$ and $e_{q}$ as defined previously. First, we obtain patch features by inputting the images into a Visual Transformer architecture. We use a frozen DINOv2 [35] backbone. These features are then fed into an encoder head, comprising one linear projection layer plus a few fully connected layers with GELU activation and a dropout layer to encode the centers of the patches. Given the corresponding patch pair $(p,q)$ as a positive sample, we train this encoder to predict similar embeddings for $p$ and $q$. To do so, we employ contrastive learning on the similarities over the trained embeddings.
@@
For each query patch $p$, we prepared a set of candidate patches $\{q,q_{1},q_{2},...,q_{n}\}$ for training, where $\{q_{1},q_{2},...,q_{n}\}$ act as $n$ negative samples, depicting patches that do not have a common field-of-view with the query patch. In our training pipeline, half of the samples have no overlap (negative samples), e.g., they depict different scenes. The other half are from the same scene as the query. In the samples, there are 30-90% negative patches, which are the ones not matchable with the query patch. We consider the loss in a patch-level manner. We train our model by optimizing the contrastive loss on the similarities of the embeddings. The contrastive loss on patch-to-patch matching pairs is defined as:
@@
$\begin{split}\mathcal{L}_{\text{contrastive}}=\frac{1}{n^{2}}\sum_{i}^{n}\sum_{j}^{n}\biggl{(}l_{\text{GT}}*(\sigma-\delta(e_{p},e_{q}))^{2}\\ -(1-l_{\text{GT}})*\delta(e_{p},e_{q})^{2}\biggr{)},\end{split}$ (6)
@@
where $l_{\text{GT}}$ is the ground-truth labels for each pair of patches (one from the query image and the other from the database). It is positive if their similarity is larger than a threshold $\llbracket sim(p_{i},q_{j})>\epsilon\rrbracket$ and vice visa. Parameter $\delta(\cdot)$ is the similarity measurement of the $d$-dimensional descriptors, i.e., cosine similarity with $\sigma=1$ as the margin. It penalizes the cases of high similarities on the patches but non-overlapping in the ground truth, i.e., no visible 3D points. We resize the input images to a fixed size. Therefore, the overall loss is averaged over $n^{2}$ tentative matching patches on one image pair. We also experimented with incorporating attentions into the procedure. However, we noticed that the attention layers only increase the size of the network while leading to worse generalization performance. Thus, we keep only the contrastive loss for training.
@@
To provide means for quick prefiltering, we add a classification ([CLS]) token to the employed Transformer network that is trained to distill global image embeddings based on the image overlap scores, shown in the ablations. These tokens are used to efficiently select a subset of potentially overlapping images given the query.
@@
Constructing Training Data. Now, we introduce the visual supervision used to build GT labels for training. The ground-truth overlapping patches of the training image pairs are achieved following LightGlue [29] employing pixel-wise depth and relative pose. As shown in Eq. 2, the patch overlap is measured by counting the co-visible 3D points in the reconstruction. The ground-truth matching local patches are found in two steps: First, a set of dense pixel coordinates $c_{i}$ are detected in the first image $I_{i}$, and transformed to 3D points $X_{i}$. The obtained 3D point cloud is filtered by non-zero depths and transformed to the reference camera, i.e., projecting the keypoints from camera $i$ to $j$ by the given transformation matrix, $kp_{i2j}$. Given camera intrinsics, we project the 3D points $X_{i}$ to the camera planes and check for visibility. We make sure that only cycle-consistent points are kept by using the projected point, calculating its 3D coordinates, and back-projecting to the original image for a visibility check again. The ground-truth matches are supposed to contain the points with a consistent depth [29]. This procedure provides the initial set of positive and negative pairs of matching image patches.
@@
Next, given the visible pixel matches, we check which patches they belong to and define the matchable patches by having at least one visible pixel correspondence. We assume one-to-one matching relations, i.e., the patch with the most patch matches in the image pair will be chosen if there are multiple. Note that the visual supervision is built mutually from the image to the reference and vice versa.